# High Resolution UDF Meshing via Iterative Networks

### v2

@@ A.2 Implementation details and other ablation studies
MGN* [Bhatnagar19] ShapeNet cars [Chang15] ShapeNet chairs [Chang15] ShapeNet planes [Chang15]
@@
128-DC UNDC [Chen22b] 0.651 6.80 1.43 11.6 12.0 3.26 94.4 15.5 11.2 3.33 5.77 2.69
DualMesh-UDF [Zhang23b] 2470 25.6 25.4 2090 12.0 13.2 12800 12.0 11.3 995 10.8 12.2
DualMesh-UDF-T [Zhang23b] 0.561 6.51 1.36 4.17 10.5 2.10 29.3 13.5 5.94 2.58 5.34 1.94
NSD-UDF + DualMesh-UDF [Stella24] 0.546 6.48 1.50 5.14 11.4 2.21 27.2 13.3 5.55 2.76 4.96 2.33
@@
256-DC UNDC [Chen22b] 1.25 6.80 2.41 97.7 11.9 9.59 716 17.6 16.1 21.1 6.89 5.63
DualMesh-UDF [Zhang23b] 2280 25.9 25.5 1390 11.6 12.8 12900 11.9 10.9 241 9.85 11.8
DualMesh-UDF-T [Zhang23b] 0.673 6.49 1.31 9.94 10.7 3.12 72.3 13.6 8.97 2.96 3.96 2.39
NSD-UDF + DualMesh-UDF [Stella24] 0.565 6.80 1.37 11.1 11.8 3.10 71.2 13.1 8.33 3.07 4.44 3.16
@@
512-DC UNDC [Chen22b] 11.3 9.36 6.95 - - - - - - - - -
DualMesh-UDF [Zhang23b] 2080 26.0 25.6 1960 11.4 12.5 14500 11.6 10.3 267 9.88 11.7
DualMesh-UDF-T [Zhang23b] 1.36 7.14 2.51 38.5 11.5 5.96 221 13.6 14.3 6.87 4.94 3.93
NSD-UDF + DualMesh-UDF [Stella24] 1.18 7.48 2.92 69.7 12.6 7.41 467 14.9 15.5 11.4 5.2 4.83
@@
CAP-L scenes [Zhou24] CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]
@@
128-MC CAP-UDF [Zhou24] 4.27 68.0 83.8 11.4 48.5 85.3 7.96 60.1 86.2
MeshUDF [Guillard22b] 4.40 68.2 84.5 11.9 49.7 86.1 9.52 61.4 87.0
DCUDF-T [Hou23a] 279 59.7 84.0 131 45.0 83.7 99.4 53.2 87.0
DCUDF-T-nocut [Hou23a] - - - 10.7 51.3 87.6 15.3 58.6 90.2
NSD-UDF + MC [Stella24] 3.34 69.8 86.7 11.1 51.7 86.3 4.86 66.0 87.8
@@
256-MC CAP-UDF [Zhou24] 3.65 70.1 86.1 10.1 53.4 86.4 4.04 69.0 86.6
MeshUDF [Guillard22b] 3.42 69.6 86.3 10.2 52.3 86.1 5.61 66.8 87.5
DCUDF-T [Hou23a] 4.75 69.9 84.6 42.1 50.7 84.7 106 63.8 84.6
DCUDF-T-nocut [Hou23a] - - - 12.0 53.1 85.7 6.37 68.8 87.1
NSD-UDF + MC [Stella24] 3.30 70.3 86.3 11.5 52.9 86.1 3.85 70.6 86.7
@@
512-MC CAP-UDF [Zhou24] 3.57 70.5 86.1 10.8 53.8 86.2 4.87 69.3 84.0
MeshUDF [Guillard22b] 4.49 69.9 84.5 10.9 52.7 85.7 4.25 68.8 86.2
DCUDF-T [Hou23a] 140 68.9 83.3 37.0 50.9 85.0 727 61.8 78.3
DCUDF-T-nocut [Hou23a] - - - 13.5 51.5 84.4 6.64 68.2 86.4
NSD-UDF + MC [Stella24] 3.72 70.0 84.1 12.3 52.5 85.6 5.92 67.1 81.1
@@
CAP-L scenes [Zhou24] CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]
@@
128-MC CAP-UDF [Zhou24] 1.98 19.9 2.43 4.71 12.2 3.07 3.88 11.2 2.53
MeshUDF [Guillard22b] 1.38 20.1 1.42 5.77 13.9 3.13 5.23 13.4 2.97
DCUDF-T [Hou23a] 385 20.1 6.53 318 12.9 5.91 109 12.0 5.37
DCUDF-T-nocut [Hou23a] - - - 5.37 12.6 3.06 12.5 12.5 1.65
NSD-UDF + MC [Stella24] 1.67 20.8 2.20 6.91 14.2 3.31 3.54 12.6 2.69
@@
256-MC CAP-UDF [Zhou24] 1.91 21.1 1.94 6.65 14.4 3.46 2.35 12.6 3.41
MeshUDF [Guillard22b] 1.79 21.3 2.03 5.65 14.9 3.48 3.06 13.3 2.78
DCUDF-T [Hou23a] 3.71 21.8 2.99 46.7 15.5 4.38 202 14.3 7.80
DCUDF-T-nocut [Hou23a] - - - 6.85 15.2 4.13 3.66 13.4 3.40
NSD-UDF + MC [Stella24] 1.76 21.5 2.63 7.27 15.1 3.69 2.47 13.0 3.56
@@
512-MC CAP-UDF [Zhou24] 1.98 21.3 2.44 7.26 15 3.64 3.14 14.2 5.29
MeshUDF [Guillard22b] 3.38 21.5 3.45 6.35 15.3 3.94 2.46 13.3 3.30
DCUDF-T [Hou23a] 191 21.3 4.07 28.2 15.3 4.21 1310 15.8 12.2
DCUDF-T-nocut [Hou23a] - - - 7.45 15.3 5.10 3.78 13.5 3.57
NSD-UDF + MC [Stella24] 2.07 21.1 3.21 7.66 15.4 4.07 3.95 15.6 7.05
@@
CAP-L scenes [Zhou24] CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]
@@
128-DC UNDC [Chen22b] 3.34 69.7 84.4 12.7 52.3 84.8 4.76 60.9 87.6
DualMesh-UDF [Zhang23b] 53.5 63.4 69.5 11.3 54.0 85.0 5.77 62.5 84.3
NSD-UDF + DualMesh-UDF [Stella24] 33.6 69.8 84.5 7.01 56.0 87.0 3.51 68.3 88.3
@@
256-DC UNDC [Chen22b] 7.85 69.1 76.9 9.89 52.9 83.9 17.5 58.3 81.2
DualMesh-UDF [Zhang23b] 50.4 63.0 68.3 10.8 53.2 84.3 5.43 61.1 82.5
NSD-UDF + DualMesh-UDF [Stella24] 34.3 68.9 83.4 8.42 54.0 86.2 3.08 71.5 87.7
@@
512-DC UNDC [Chen22b] - - - - - - - - -
DualMesh-UDF [Zhang23b] 51.7 59.8 61.8 11.4 50.6 83.1 5.69 58.2 79.0
NSD-UDF + DualMesh-UDF [Stella24] 35.2 66.5 77.9 10.1 51.7 85.4 3.99 68.8 85.1
@@
CAP-L scenes [Zhou24] CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]
@@
128-DC UNDC [Chen22b] 2.55 20.2 3.97 13.4 13.0 3.27 2.22 13.1 2.14
DualMesh-UDF [Zhang23b] 35.7 18.9 7.07 6.58 13.8 4.27 2.77 13.7 3.42
NSD-UDF + DualMesh-UDF [Stella24] 42.8 17.6 0.739 3.98 13.1 2.88 2.04 11.7 2.14
@@
256-DC UNDC [Chen22b] 8.26 21.8 9.61 5.38 14.4 3.88 30.5 10.3 1.65
DualMesh-UDF [Zhang23b] 35.7 19.5 7.01 6.34 14.7 4.60 2.66 13.9 3.16
NSD-UDF + DualMesh-UDF [Stella24] 43.3 18.6 0.403 4.82 14.3 3.51 1.96 12.5 2.82
@@
512-DC UNDC [Chen22b] - - - - - - - - -
DualMesh-UDF [Zhang23b] 35.0 19.9 8.32 6.45 14.6 4.86 2.68 13.5 3.03
NSD-UDF + DualMesh-UDF [Stella24] 42.3 18.6 2.04 5.74 14.7 4.04 2.50 14.0 3.80
@@ A.3 Training convergence
In the method section of the main paper we state that our pipeline applies a sigmoid function to the network outputs before using them as input for the next iteration. We have experimentally found that, when using an identity activation (i.e. direct input) and training with only one iteration, the network goes from an IC of 87.0 on ShapeNet [Chang15] cars at resolution 512 to 88.6, however it starts diverging after that. Using a softmax activation, which helps normalizing the otherwise unbound network outputs, the training converges, but the network does not produce significant improvements over iterations, going from an IC of 87.3 to 87.5 after two iterations. Using a sigmoid activation, instead, showed a more steady improvement over iterations, going from 87.6 to 88.5 after one iteration, and 88.9 after an additional one. However, training it with more than one iteration did not show significant improvements. Using a random number of training iterations, instead, helped the network to converge until 5 iterations, achieving similar IC scores but better CD scores ($9.90\times 10^{-5}$ vs $8.84\times 10^{-5}$), as shown in the main paper, signifying an overall similar accuracy but better surface retrieval capabilities at high resolutions.
@@ A.4 Auto-decoder training
For our auto-decoder experiments we used the traditional auto-decoder architecture proposed in DeepSDF [Park19c]. The input meshes are rescaled and centered within a $1,1]^{3}$ volume, and during the data preparation phase, training points are sampled. For each mesh, 20k points are uniformly sampled within the volume, while 400k points are sampled near the surface. To obtain the surface points, 200k points are first uniformly distributed on the surface, and then small amounts of Gaussian noise are added. Gaussian noise with a mean of 0 and a standard deviation of $\sqrt{0.005}$ is applied to the first 200k surface points, and noise with a mean of 0 and a standard deviation of $\sqrt{0.0005}$ is added to the remaining 200k points. The auto-decoder network consists of 12 layers, each with 1024 hidden nodes and ReLU activations, and latent codes of size 512. It is trained using L1 loss, without regularization or Fourier encoding, for 10k epochs with a batch size of 16. To focus the network’s capacity on the surface, the UDF is clamped to 0.1. The Adam optimizer [Kingma15a] is used with learning rates of 0.0005 for the model and 0.001 for the latent codes, with learning rate decay applied at epochs 1600 and 3500 by a factor of 0.35.
@@ A.5 Different auto-decoder UDF architecture
Table A.7: Triangulating a Softplus-based auto-decoder. L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions on the ShapeNet [Chang15] cars dataset. The best results are in bold.
@@
To test the robustness of our method to different UDF architectures, we trained an auto-decoder with a different architecture: using a softplus activation function and Eikonal loss with a weight of 0.1, with the rest as in the section above. We show the results on ShapeNet [Chang15] cars in Tab. A.7, along with the highest-scoring baselines from the main experiment. While all methods achieved slightly better performance compared to the ReLU-based architecture, the same conclusions apply. Our method outperforms the tested baselines, particularly so at high resolutions, while also achieving lower standard deviations across the dataset.
@@ A.6 Additional figures
We show here additional qualitative results of our method compared to the baselines. In Fig.˜A.1 we show an additional example at different resolutions compared to the NSD-UDF [Stella24] baseline: the shapes look similar at low resolutions, but at 256 and 512 the baseline cannot retrieve large portions of the surface.
@@
In Fig. A.7, we show the meshing on the CAP-L 3D scene [Zhou24] used in Fig. 4 of the main text in greater size to better appreciate the details, e.g., the base of the statues.
@@
Figure A.1: Meshing at different resolutions, additional examples. While NSD-UDF [Stella24] retrieves most of the surface well at a low resolution, it struggles at higher ones. In contrast, our method, recovers the surface well at all resolutions. We use Marching Cubes with both methods.
@@
Figure A.7: Meshing of a 3D scene from CAP-L [Zhou24]. Surface meshing results of the "Burghers" scene with all methods at resolution of 512. UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is Dual Contouring-based, otherwise it is Marching Cubes-based.
A.7 DoubleCoverUDF [Hou23a] tuning and min-cut
To achieve better results with DCUDF [Hou23a] compared to its default parameters, we have run the algorithm 3 times per experiment and per resolution, with different parameters, and we have selected the best run in each scenario. Additionally, in Fig. A.8, we present the meshing results of DCUDF [Hou23a] on UDFs from different models on cars, both with and without the final min-cut step, at resolutions of 128 and 512. Without the min-cut, more surfaces are retained, though they are double-layered. Moreover, the overall meshing quality deteriorates at higher resolutions, following the trend observed for the other baselines.
Auto-decoder cars CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]

### v3

@@ A.2 Implementation details and other ablation studies
MGN* [1] ShapeNet cars [3] ShapeNet chairs [3] ShapeNet planes [3]
@@
128-DC UNDC [5] 0.651 6.80 1.43 11.6 12.0 3.26 94.4 15.5 11.2 3.33 5.77 2.69
DualMesh-UDF [41] 2470 25.6 25.4 2090 12.0 13.2 12800 12.0 11.3 995 10.8 12.2
DualMesh-UDF-T [41] 0.561 6.51 1.36 4.17 10.5 2.10 29.3 13.5 5.94 2.58 5.34 1.94
NSD-UDF + DualMesh-UDF [36] 0.546 6.48 1.50 5.14 11.4 2.21 27.2 13.3 5.55 2.76 4.96 2.33
@@
256-DC UNDC [5] 1.25 6.80 2.41 97.7 11.9 9.59 716 17.6 16.1 21.1 6.89 5.63
DualMesh-UDF [41] 2280 25.9 25.5 1390 11.6 12.8 12900 11.9 10.9 241 9.85 11.8
DualMesh-UDF-T [41] 0.673 6.49 1.31 9.94 10.7 3.12 72.3 13.6 8.97 2.96 3.96 2.39
NSD-UDF + DualMesh-UDF [36] 0.565 6.80 1.37 11.1 11.8 3.10 71.2 13.1 8.33 3.07 4.44 3.16
@@
512-DC UNDC [5] 11.3 9.36 6.95 - - - - - - - - -
DualMesh-UDF [41] 2080 26.0 25.6 1960 11.4 12.5 14500 11.6 10.3 267 9.88 11.7
DualMesh-UDF-T [41] 1.36 7.14 2.51 38.5 11.5 5.96 221 13.6 14.3 6.87 4.94 3.93
NSD-UDF + DualMesh-UDF [36] 1.18 7.48 2.92 69.7 12.6 7.41 467 14.9 15.5 11.4 5.2 4.83
@@
CAP-L scenes [46] CAP-L cars [46] DiffUDF cars [11]
@@
128-MC CAP-UDF [46] 4.27 68.0 83.8 11.4 48.5 85.3 7.96 60.1 86.2
MeshUDF [13] 4.40 68.2 84.5 11.9 49.7 86.1 9.52 61.4 87.0
DCUDF-T [14] 279 59.7 84.0 131 45.0 83.7 99.4 53.2 87.0
DCUDF-T-nocut [14] - - - 10.7 51.3 87.6 15.3 58.6 90.2
NSD-UDF + MC [36] 3.34 69.8 86.7 11.1 51.7 86.3 4.86 66.0 87.8
@@
256-MC CAP-UDF [46] 3.65 70.1 86.1 10.1 53.4 86.4 4.04 69.0 86.6
MeshUDF [13] 3.42 69.6 86.3 10.2 52.3 86.1 5.61 66.8 87.5
DCUDF-T [14] 4.75 69.9 84.6 42.1 50.7 84.7 106 63.8 84.6
DCUDF-T-nocut [14] - - - 12.0 53.1 85.7 6.37 68.8 87.1
NSD-UDF + MC [36] 3.30 70.3 86.3 11.5 52.9 86.1 3.85 70.6 86.7
@@
512-MC CAP-UDF [46] 3.57 70.5 86.1 10.8 53.8 86.2 4.87 69.3 84.0
MeshUDF [13] 4.49 69.9 84.5 10.9 52.7 85.7 4.25 68.8 86.2
DCUDF-T [14] 140 68.9 83.3 37.0 50.9 85.0 727 61.8 78.3
DCUDF-T-nocut [14] - - - 13.5 51.5 84.4 6.64 68.2 86.4
NSD-UDF + MC [36] 3.72 70.0 84.1 12.3 52.5 85.6 5.92 67.1 81.1
@@
CAP-L scenes [46] CAP-L cars [46] DiffUDF cars [11]
@@
128-MC CAP-UDF [46] 1.98 19.9 2.43 4.71 12.2 3.07 3.88 11.2 2.53
MeshUDF [13] 1.38 20.1 1.42 5.77 13.9 3.13 5.23 13.4 2.97
DCUDF-T [14] 385 20.1 6.53 318 12.9 5.91 109 12.0 5.37
DCUDF-T-nocut [14] - - - 5.37 12.6 3.06 12.5 12.5 1.65
NSD-UDF + MC [36] 1.67 20.8 2.20 6.91 14.2 3.31 3.54 12.6 2.69
@@
256-MC CAP-UDF [46] 1.91 21.1 1.94 6.65 14.4 3.46 2.35 12.6 3.41
MeshUDF [13] 1.79 21.3 2.03 5.65 14.9 3.48 3.06 13.3 2.78
DCUDF-T [14] 3.71 21.8 2.99 46.7 15.5 4.38 202 14.3 7.80
DCUDF-T-nocut [14] - - - 6.85 15.2 4.13 3.66 13.4 3.40
NSD-UDF + MC [36] 1.76 21.5 2.63 7.27 15.1 3.69 2.47 13.0 3.56
@@
512-MC CAP-UDF [46] 1.98 21.3 2.44 7.26 15 3.64 3.14 14.2 5.29
MeshUDF [13] 3.38 21.5 3.45 6.35 15.3 3.94 2.46 13.3 3.30
DCUDF-T [14] 191 21.3 4.07 28.2 15.3 4.21 1310 15.8 12.2
DCUDF-T-nocut [14] - - - 7.45 15.3 5.10 3.78 13.5 3.57
NSD-UDF + MC [36] 2.07 21.1 3.21 7.66 15.4 4.07 3.95 15.6 7.05
@@
CAP-L scenes [46] CAP-L cars [46] DiffUDF cars [11]
@@
128-DC UNDC [5] 3.34 69.7 84.4 12.7 52.3 84.8 4.76 60.9 87.6
DualMesh-UDF [41] 53.5 63.4 69.5 11.3 54.0 85.0 5.77 62.5 84.3
NSD-UDF + DualMesh-UDF [36] 33.6 69.8 84.5 7.01 56.0 87.0 3.51 68.3 88.3
@@
256-DC UNDC [5] 7.85 69.1 76.9 9.89 52.9 83.9 17.5 58.3 81.2
DualMesh-UDF [41] 50.4 63.0 68.3 10.8 53.2 84.3 5.43 61.1 82.5
NSD-UDF + DualMesh-UDF [36] 34.3 68.9 83.4 8.42 54.0 86.2 3.08 71.5 87.7
@@
512-DC UNDC [5] - - - - - - - - -
DualMesh-UDF [41] 51.7 59.8 61.8 11.4 50.6 83.1 5.69 58.2 79.0
NSD-UDF + DualMesh-UDF [36] 35.2 66.5 77.9 10.1 51.7 85.4 3.99 68.8 85.1
@@
CAP-L scenes [46] CAP-L cars [46] DiffUDF cars [11]
@@
128-DC UNDC [5] 2.55 20.2 3.97 13.4 13.0 3.27 2.22 13.1 2.14
DualMesh-UDF [41] 35.7 18.9 7.07 6.58 13.8 4.27 2.77 13.7 3.42
NSD-UDF + DualMesh-UDF [36] 42.8 17.6 0.739 3.98 13.1 2.88 2.04 11.7 2.14
@@
256-DC UNDC [5] 8.26 21.8 9.61 5.38 14.4 3.88 30.5 10.3 1.65
DualMesh-UDF [41] 35.7 19.5 7.01 6.34 14.7 4.60 2.66 13.9 3.16
NSD-UDF + DualMesh-UDF [36] 43.3 18.6 0.403 4.82 14.3 3.51 1.96 12.5 2.82
@@
512-DC UNDC [5] - - - - - - - - -
DualMesh-UDF [41] 35.0 19.9 8.32 6.45 14.6 4.86 2.68 13.5 3.03
NSD-UDF + DualMesh-UDF [36] 42.3 18.6 2.04 5.74 14.7 4.04 2.50 14.0 3.80
@@ A.3 Training convergence
In the method section of the main paper we state that our pipeline applies a sigmoid function to the network outputs before using them as input for the next iteration. We have experimentally found that, when using an identity activation (i.e. direct input) and training with only one iteration, the network goes from an IC of 87.0 on ShapeNet [3] cars at resolution 512 to 88.6, however it starts diverging after that. Using a softmax activation, which helps normalizing the otherwise unbound network outputs, the training converges, but the network does not produce significant improvements over iterations, going from an IC of 87.3 to 87.5 after two iterations. Using a sigmoid activation, instead, showed a more steady improvement over iterations, going from 87.6 to 88.5 after one iteration, and 88.9 after an additional one. However, training it with more than one iteration did not show significant improvements. Using a random number of training iterations, instead, helped the network to converge until 5 iterations, achieving similar IC scores but better CD scores ($9.90\times 10^{-5}$ vs $8.84\times 10^{-5}$), as shown in the main paper, signifying an overall similar accuracy but better surface retrieval capabilities at high resolutions.
@@ A.4 Auto-decoder training
For our auto-decoder experiments we used the traditional auto-decoder architecture proposed in DeepSDF [27]. The input meshes are rescaled and centered within a $1,1]^{3}$ volume, and during the data preparation phase, training points are sampled. For each mesh, 20k points are uniformly sampled within the volume, while 400k points are sampled near the surface. To obtain the surface points, 200k points are first uniformly distributed on the surface, and then small amounts of Gaussian noise are added. Gaussian noise with a mean of 0 and a standard deviation of $\sqrt{0.005}$ is applied to the first 200k surface points, and noise with a mean of 0 and a standard deviation of $\sqrt{0.0005}$ is added to the remaining 200k points. The auto-decoder network consists of 12 layers, each with 1024 hidden nodes and ReLU activations, and latent codes of size 512. It is trained using L1 loss, without regularization or Fourier encoding, for 10k epochs with a batch size of 16. To focus the network’s capacity on the surface, the UDF is clamped to 0.1. The Adam optimizer [16] is used with learning rates of 0.0005 for the model and 0.001 for the latent codes, with learning rate decay applied at epochs 1600 and 3500 by a factor of 0.35.
@@ A.5 Different auto-decoder UDF architecture
Table A.7: Triangulating a Softplus-based auto-decoder. L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions on the ShapeNet [3] cars dataset. The best results are in bold.
@@
To test the robustness of our method to different UDF architectures, we trained an auto-decoder with a different architecture: using a softplus activation function and Eikonal loss with a weight of 0.1, with the rest as in the section above. We show the results on ShapeNet [3] cars in Tab. A.7, along with the highest-scoring baselines from the main experiment. While all methods achieved slightly better performance compared to the ReLU-based architecture, the same conclusions apply. Our method outperforms the tested baselines, particularly so at high resolutions, while also achieving lower standard deviations across the dataset.
@@ A.6 Additional figures
We show here additional qualitative results of our method compared to the baselines. In Fig.˜A.1 we show an additional example at different resolutions compared to the NSD-UDF [36] baseline: the shapes look similar at low resolutions, but at 256 and 512 the baseline cannot retrieve large portions of the surface.
@@
In Fig. A.7, we show the meshing on the CAP-L 3D scene [46] used in Fig. 4 of the main text in greater size to better appreciate the details, e.g., the base of the statues.
@@
Figure A.1: Meshing at different resolutions, additional examples. While NSD-UDF [36] retrieves most of the surface well at a low resolution, it struggles at higher ones. In contrast, our method, recovers the surface well at all resolutions. We use Marching Cubes with both methods.
@@
Figure A.7: Meshing of a 3D scene from CAP-L [46]. Surface meshing results of the "Burghers" scene with all methods at resolution of 512. UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is Dual Contouring-based, otherwise it is Marching Cubes-based.
A.7 DoubleCoverUDF [14] tuning and min-cut
To achieve better results with DCUDF [14] compared to its default parameters, we have run the algorithm 3 times per experiment and per resolution, with different parameters, and we have selected the best run in each scenario. Additionally, in Fig. A.8, we present the meshing results of DCUDF [14] on UDFs from different models on cars, both with and without the final min-cut step, at resolutions of 128 and 512. Without the min-cut, more surfaces are retained, though they are double-layered. Moreover, the overall meshing quality deteriorates at higher resolutions, following the trend observed for the other baselines.
Auto-decoder cars CAP-L cars [46] DiffUDF cars [11]